OpenStudy (anonymous):

graph y = x-3 0 4, y = x - 3 - 4, y = -2 x+ 3 - 4 and y = 2 x + 3 - 4

OpenStudy (anonymous):

is that.. y=x-304 or there's a multiply between each number like.. y=x-3*0*4 ??

OpenStudy (anonymous):

no its not . the zero wasnt ment to be there

OpenStudy (anonymous):

its x-3 - 4,y

OpenStudy (anonymous):

graph y = x-3 - 4, y = x - 3 - 4, y = -2 x+ 3 - 4 and y = 2 x + 3 - 4

OpenStudy (anonymous):

... here you are !